OpenStudy (cookiemonster18):

(PLEASE CORRECT MY ANSWERS IF IM WRONG,I'LL GIVE MEDALS) 1.Write an equation in slope- intercept form of the line that passes through the given point and is parallel to the graph of the given equation. (2,-2);y=-x-2 A.y=-2x*** B.y=2x C.y=1/2x D.y=-x 2.Write an equation in slope- intercept form of the line that passes through the given point and is parallel to the graph of the given equation. (2,-1);y=-3/2x-6 A.y=-3/2x+1 B.y=-3/2x-1*** C.y=-3/2x+2 D.y=-3/2x+4

OpenStudy (cookiemonster18):

3.Write an equation in slope- intercept form of the line that passes through the given point and is parallel to the graph of the given equation. (4,2);x=-3 A.y=2 B.y=2x+4*** C.y=4x D.y=4 4.Write an equation in slope- intercept form of the line that passes through the given point and is perpendicular to the graph of the given equation. (-2,3);y=1/2x-1 A.y=1/2x+1 B.y=-2x-1*** C.y=1/2x-1 D.y=-1/2x-1

OpenStudy (cookiemonster18):

5.Write an equation in slope- intercept form of the line that passes through the given point and is perpendicular to the graph of the given equation. (5,0);y+1=2(x-3) A.y=-1/2x+5 B.y=2x-5 C.y=1/2x-2*** D.y=-1/2x+5/2

OpenStudy (cookiemonster18):

@Vocaloid

OpenStudy (anonymous):

what your answer?

OpenStudy (cookiemonster18):

the ones with ----> *** are my answers @ArifL

OpenStudy (dayakar):

1) is wrong

OpenStudy (dayakar):

can u your work ,i help u where u make wrong

OpenStudy (dayakar):

can show ur work

OpenStudy (dayakar):

2) is also wrong

OpenStudy (dayakar):

1)D 2)C

OpenStudy (cookiemonster18):

what about 3,4, and 5? are those right or wrong too? @dayakar

OpenStudy (dayakar):

4)correct

OpenStudy (dayakar):

5)A

OpenStudy (dayakar):

3) verify the problem there is line which is parallel x=-3 and passes through given point

OpenStudy (dayakar):

there is no such line